tdfygukhikjhbgfch:

Nova receives a bill for her cell phone each month. She pays a flat fee of $35 plus $0.05 per text message. This month, her bill without taxes was $51.50. Which equation gives the number of text messages, t, she sent this month?

tdfygukhikjhbgfch:

please help

tdfygukhikjhbgfch:

i dont understand

Hoodmemes:

Ight do \[35-51.50 \] divide that by 0.5 and u should egt ya answer bro

tdfygukhikjhbgfch:

A. 35+0.05⋅t=51.50 b. (35+0.05)t=51.50 c. 35+0.05=51.50⋅t d. 35⋅t+0.05=51.50

tdfygukhikjhbgfch:

;(

tdfygukhikjhbgfch:

bye

jhonyy9:

@Hoodmemes i think may be more logical subtracting 35 from 51,50 and so in this way get a positiv result please check your calcule

jhonyy9:

@Vocaloid

Hoodmemes:

Oh so if the cost of per message is 0.05 and the cost of them text messages be 0.05,and she be paying a flat fee of like 35,and her total bill without taxes be 51.50,then i must be having to use addition but also multiplication to solve this problem,and That would leave two answer choices left A and D?